consider a market with inverse demand curve p=400-4Q.

costs per firm are given by C(q)= 16+10q+q^2

a) find the minimum efficient scale output level

b) in a competitive market , how many firms will be active in the long-run

c) suppose we have a cournot oligopoly with n firms . determine output of each firm and the equilibrium price

d) find the long run equilibrium number of firms if the market is a cournot oligopoly and entry occurs until profit per firm is zero
Answer #1

2 4 ) P=400-40, 6(9)= 16 +109 +92090 a) at MES, AG is Minimum AC=C/9 = 16/9 + 10 + q dace = -16.2 +4 =0, 9-4. Ang b.) In long

c.) Cournot game, n firms, Q = + 2q = 4 + Q - V then To= pan-cow) = (400 - 494 - 48-i) W - 16-10 -4² I = (400 - 4Q i) W - 40080 P = 400 4x39 (06 to 40) - 400 - 1560 To6 to 40) d.) LR Eq m * 8o when profit per from is zero, n î To= pq - Cac) P. - (940